A particle f mass 5kg is moving in anticlockwise direction with aspect of 3m/s ina circular path of radius 10cm.Find the angle moving of this particle with respect to the centre of the path?


1
Expert's answer
2021-02-28T07:31:53-0500

"v=\omega*r \text{ where }\omega-\text{angular velocity};"

"\omega= \frac{v}{r}=\frac{3}{0.1}=30"

"\theta(t)= \omega*t = 30t\text{ where }\theta- \text{counterclockwise rotation angle}"

"\text{or }\theta(t) = -30t\text{ where }\theta- \text{clockwise rotation}"